You have a choice when subscribing toour magazine:you can

(i) pay R100 now for a next four year subscription,

(ii) pay R28 at the beginning of each month for four years,or

(iii) pay R54 today and R54 again two years from today.

Which one is the best deal for you,the subscribers,if your opportunity cost of funds is 10%?



1
Expert's answer
2021-11-04T16:50:05-0400


i) Present Value of the cost of option a = $100


ii)Present Value of the cost of option b "= \$28+\frac{28}{(1.1)^1} +\frac{18}{(1.1)^2} +\frac{28}{(1.1)^3} = \$97.63"

iii)Present Value = "\$54 +\frac{\$54}{(1+0.10)^2}"

"=\$54+44.63 = \$98.63"


the best deal is option (ii),it is the lowest. Pay R28 at the beginning of each month for four years